# Robust Design and Evaluation of

### v5

@@ 6.1 Bounding individual default risk
We first use our nonparametric regression procedures in Section 3 to construct alternative confounding-robust credit risk scores and explore how varying assumptions of unobserved confounding affects predictions of individual default risk.
@@
We estimate the upper bound on the conditional default probability with nonparametric outcome bounds, $\overline{\mu}(\cdot)$. We split our sample of applications into two folds, using the first fold to estimate $\pi_{1}(\cdot)$, $\mu_{1}(\cdot)$ using random forests and using cross-validated logistic regression with a Lasso penalty to regress the estimated pseudo-outcome $\phi_{1}(Y_{i};\hat{\eta})+\overline{\Gamma}\phi(\pi_{0}(X_{i})\mu_{1}(X_{i});\hat{\eta})$ on application-level covariates $X_{i}$. Throughout, we set $\underline{\Gamma}=1$ and vary $\overline{\Gamma}\in\{1,2,3\}$.
@@
[…]
Notes: The left panel summarizes the joint distribution of a benchmark credit risk score’s predictions of default risk against our robust predictions of the default risk. Among applications in each decile of the benchmark credit score’s predicted risk distribution, the left panel plots the percentage of applications at each decile of our robust estimator’s predicted risk distribution. Our robust estimator is constructed assuming $\underline{\Gamma}=1$, $\overline{\Gamma}=2$, and the benchmark credit score predicts default risk among only funded applications. The right panel summarizes how the coefficients on a subset of application-level characteristics vary with our assumptions on unobserved confounding over $\overline{\Gamma}\in\{1,2,3\}$. The value $\overline{\Gamma}=1$ corresponds to the benchmark credit score.
We compare our robust estimator’s predictions of default risk against a benchmark credit score that predicts default risk among funded applications. The left panel of Figure 3b summarizes the joint distribution of the benchmark credit score’s predictions and our estimator’s predictions, where our estimator is constructed assuming $\underline{\Gamma}=1,\overline{\Gamma}=2$. Among applications in each decile of the benchmark credit score’s predicted risk distribution, the left panel plots the percentage of applications at each decile of our estimator’s predicted risk distribution. Among applications at any particular decile of the benchmark credit score’s predicted risk distribution, there exists meaningful variation in our estimator’s risk predictions. Among applications in the 3rd decile and 8th decile of the benchmark risk score’s predicted risk distribution, for example, 13.8% of applications are in the top half of our estimator’s predicted risk distribution and 4.7% of applications are in the bottom half of our estimator’s predicted risk distribution respectively. Among applications in the 5th decile of the benchmark risk score’s predicted risk distribution, 18.7% of applications are flipped to either the bottom (deciles 1-3) or top (deciles 8-10) of our estimator’s predicted risk distribution. Such differences meaningfully affect who would receive access to credit. If, for example, only the bottom third of least risky applications are funded, then 10.1% of all applications would have their decision flipped between the benchmark risk score and the robust risk score that accounts for possible unobserved confounding.
The right panel of Figure 3b investigates these differences in predicted default risk by comparing how the estimated coefficients on a subset of characteristics vary as $\overline{\Gamma}$ varies. Online Appendix Table 3 provides a more detailed description of the variable names in the right panel. Altering our assumptions on unobserved confounding leads to effectively no difference in the estimated coefficients for some notable characteristics. The estimated coefficient on the application’s total net income and credit bureau score do not vary as $\overline{\Gamma}$ varies and are always equal to zero. In contrast, for some characteristics such as the number of credit card applications submitted by all applicants on the application, the benchmark risk score places no weight them, whereas all estimators with $\overline{\Gamma}>1$ incorporate them into the model and assign a non-zero weight. For other characteristics like the applicant’s occupation type or their maximum delinquency over the last 12 months, varying our assumptions on unobserved confounding lead to large changes in the magnitudes of the estimated coefficients. Explicitly accounting for unobserved confounding leads to substantive differences in the resulting credit risk score.
6.2 Bounding the predictive performance of a credit risk score
We next apply our framework to robustly evaluate the predictive performance of an existing credit risk score. We construct a naive benchmark credit risk score by splitting our sample of personal loan applications into two subsets, which we will refer to as the training data and evaluation data. In the training data, we construct a credit risk score that predicts 5 month default risk among funded applications $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ using a random forest. In the evaluation data, we analyze this benchmark credit risk score’s mean square error and ROC curve, following recent empirical analyses of credit risk scores in mortgage lending in BlattnerNelson(21); FusterEtAl(22)predictably. We specify nonparametric outcome bounds, which restrict that the unobserved default rate among unfunded applications cannot be too different than the observed 5 month default rate among funded applicants conditional on covariates. We evaluate the benchmark credit risk score as we vary the choice of $\overline{\Gamma}$, fixing $\underline{\Gamma}=1$ throughout.
@@ 6.2 Bounding the predictive performance of a credit risk score
[…]
Notes: This figure summarizes how the bounds on the benchmark risk score’s mean square error and ROC curve vary with our assumptions on unobserved confounding. In Panel (A), the mean square error among only funded applications is plotted in the red dashed line. The bounds (black) are constructed using nonparametric outcome bounds, varying $\overline{\Gamma}\in[0,1]$ and setting $\underline{\Gamma}=1$. In Panel (B), the ROC curve among only funded applications is plotted in black. The bounds on the ROC curve for alternative choices $\overline{\Gamma}\in\{1.25,1.33,1.5\}$ are plotted in different colors.
Figure 4b(a) reports how the estimated upper and lower bounds (black) on the benchmark credit risk score’s mean square error varies as our assumptions on unobserved confounding $\overline{\Gamma}\in[1,2]$. When $\overline{\Gamma}=1$, nonparametric outcome bounds are equivalent to assuming that the funding decision is unconfounded; therefore, the upper and lower bounds both equal the benchmark credit risk score’s mean square error among only funded applications (dashed red line; 0.143). As $\overline{\Gamma}$ increases, nonparametric outcome bounds now allows default rates among unfunded applications to be larger than that of funded applications, and so the estimated bounds widen. Nonetheless, the estimated bounds remain quite informative. At $\overline{\Gamma}=2$, which allows unfunded applications to be at most 2 times as likely to default as funded applications conditional on observed application-level features, the lower bound is 0.138 and the upper bound is 0.182.
Figure 4b(b) analogously examines how the estimated ROC curve of the benchmark credit score varies as $\overline{\Gamma}$ varies on a subsample of personal loan applications in the evaluation data. Recall from Section 2.1, the ROC curve for the benchmark risk score summarizes how the true positive rate and false positive rate of the threshold classifier $1\{s(X_{i})\geq\tau\}$ varies as the threshold $\tau$ varies over $[0,1]$. In empirical applications, ROC curves are typically summarized by the Area Under the Curve (AUC), which is a popular, scalar measure for summarizing differences in predictive power across ROC curves. For each ROC curve in Figure 4b(b), we compute the implied AUC using the trapezoidal rule. When $\overline{\Gamma}=1$ and we assume the funding decision is unconfounded, the benchmark credit risk score’s AUC is 0.637. At $\overline{\Gamma}=1.5$, the implied lower bound on the benchmark credit risk score’s AUC is 0.575 and its upper bound is 0.705.
We next investigate the benchmark credit risk score’s predictive disparities across income groups. We define $G_{i}\in\{0,1\}$ to be whether an application is below or above the median personal income of all submitted applications. We examine how the benchmark credit risk score’s mean square error and ROC curve vary across these income groups, thereby studying whether it is a “noisier” signal of 5 month default risk for lower income applications, as in BlattnerNelson(21); FusterEtAl(22)predictably.
@@
Notes: This figure summarizes how the bounds on the benchmark credit risk score’s mean square error and ROC curve vary across income groups and as our assumptions on observed confounding vary. In Panel (A), the mean square among only funded applications is plotted in the red dashed line. The bounds (black) are constructed using nonparametric outcome bounds, varying $\overline{\Gamma}\in[0,1]$ and setting $\underline{\Gamma}=1$. In Panel (B), the ROC curve among only funded applications is plotted in black. The bounds on the ROC curve for alternative choices $\overline{\Gamma}\in\{1.25,1.33,1.5\}$ are plotted in different colors.
Figure 5b(a) compares the mean square error of the benchmark credit risk across across income groups. Interestingly, for $\overline{\Gamma}=1$, the mean square error of the benchmark credit risk score is significantly larger on applications below the median income (0.169) than those above it (0.136). In other words, the benchmark credit risk score’s mean square error is 24.2% larger on funded applications below the median income than those above it. This difference persists as we vary assumptions on unmeasured confounding. The upper bound on the mean square error for applications above the median income is only larger than the lower bound on the mean square error for applications below the median income for values $\overline{\Gamma}\geq 1.69$. In order to rule out this mean square error disparity across income groups, researchers would therefore need to argue that unfunded applications are at least 1.69 times as likely to default within 5 months as funded applications conditional on covariates.
Figure 5b(b) also compares the ROC curves of the benchmark credit risk score across income groups. Intriguingly, there exists effectively no disparity across income groups based on the ROC curves. When $\overline{\Gamma}=1$, the implied AUC for applications above the median income is $0.603$, whereas the implied AUC is $0.616$ for applications below the median income. Unsurprisingly, the predictive disparities based on ROC curves therefore disappears as quickly as we allow for unobserved confounding. Any disparity in terms of ROC curves is already ruled out at $\overline{\Gamma}=1.25$ as the bounds on the AUC for appplications above the median income is $[0.567,0.653]$, and it is $[0.581,0.654]$ for applications below the median income.
7 Conclusion
This paper develops general tools for the robust design and evaluation of predictive algorithms in selectively observed data under unobserved confounding. We consider a general class of restrictions on the extent to which the outcome of interest can differ on average between unselected and selected units conditional on observed covariates and identified nuisance parameters. Such restrictions capture a wide variety of popular intuitions and strategies for addressing unobserved confounding in empirical research. A large class of predictive estimands are partially identified under these restrictions, and we develop estimators for their bounds based on debiased machine learning. Ultimately, we recommend that empirical researchers report robust bounds on predicted risk and predictive performance estimands under restrictions on unobserved confounding motivated by domain-specific knowledge or empirical strategies. As predictive algorithms now inform high-stakes decisions across widely studied economic settings like consumer finance, criminal justice, health, and many others, this paper provides a tractable pipeline for robustly constructing and evaluating predictive algorithms under alternative assumptions on unobserved confounding.
Appendix A Proofs of Results in the Main Text
A.1 Proof of Lemma 2.1
@@ Appendix A Proofs of Results in the Main Text
Observe that if $\widetilde{\mbox{perf}}_{+}(s;\beta)\in\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, then there exists some $\tilde{\delta}\in\Delta$ such that $\widetilde{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\delta})$. It follows immediately that $\widetilde{\mbox{perf}}_{+}(s;\beta)\in[\underline{\mbox{perf}}_{+}(s;\beta),\overline{\mbox{perf}}_{+}(s;\beta)]$. All that remains to show is that every value in the interval $[\underline{\mbox{perf}}_{+}(s;\beta,\Delta),\overline{\mbox{perf}}_{+}(s;\beta,\Delta)]$ is achieved by some $\delta(\cdot)\in\Delta$.
Towards this, we apply a one-to-one change-of-variables. Let $U(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ be defined as $U(x)=\frac{\delta(x)-\underline{\delta}(x)}{\overline{\delta}(x)-\underline{\delta}(x)}$. For any $\delta(\cdot)\in\Delta$, there exists $U(\cdot)\in[0,1]$ such that $\mbox{perf}_{+}(s;\beta,\delta)=\mbox{perf}_{+}(s;\beta,U)$, where
@@
So, for any $U(\cdot)\in[0,1]$, there exists $\tilde{V},\tilde{U}(\cdot)$ satisfying $\tilde{U}(\cdot)\in[0,\tilde{V}]$, $\tilde{V}\geq 0$ and $\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})]\tilde{V}+\mathbb{E}[\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))\tilde{U}(X_{i})]=1$ such that $\mbox{perf}_{+}(s;\beta,U)=\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})$, where
@@
By the change-of-variables, there exists $\tilde{\underline{V}},\tilde{\underline{U}}(\cdot)$ and $\tilde{\overline{V}},\tilde{\overline{U}}(\cdot)$ such that
$\underline{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\underline{U}}(\cdot),\tilde{\underline{V}}),\mbox{ }\overline{\mbox{perf}}_{+}(s;\beta)=\mbox{perf}_{+}(s;\beta,\tilde{\overline{U}}(\cdot),\tilde{\overline{V}}).$
Therefore, $\tilde{p}=\lambda\mbox{perf}_{+}(s;\beta,\tilde{\underline{U}}(\cdot),\tilde{\underline{V}})+(1-\lambda)\mbox{perf}_{+}(s;\beta,\tilde{\overline{U}}(\cdot),\tilde{\overline{V}})$. Since $\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})$ is linear in $\tilde{U},\tilde{V}$, we also have that
$\tilde{p}=\mbox{perf}_{+}(s;\beta,\lambda\tilde{\underline{U}}+(1-\lambda)\tilde{\overline{U}},\lambda\tilde{\underline{V}}+(1-\lambda)\tilde{\overline{V}}).$
We can therefore apply the change-of-variables in the other direction to construct the corresponding $\tilde{\delta}(\cdot)\in\Delta$, which satisfies $\tilde{p}=\mbox{perf}_{+}(s;\beta,\tilde{\delta})$ by construction. $\Box$
A.2 Proof of Lemma 2.2
@@
This is minimized over $\mu^{*}(x)\in\mathcal{H}(\mu^{*}(x))$ at $\mu^{*}(x)=\underline{\mu}^{*}(x)$. The result for the lower bound immediately follows. The result for the upper bound follows by an analogous argument. $\Box$
A.3 Proof of Proposition 3.1
We prove the result for our estimator of the upper bound, and the same argument applies for our estimator of the lower bound. Let $\mathcal{O}_{1}$ denote the observations in the first fold and $\mathcal{O}_{2}$ denote the observations in the second fold. We first observe that
@@
Under Assumption B.1, $\|\widehat{\overline{\mu}}(\cdot)-\widehat{\overline{\mu}}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma B.1. Furthermore, $\hat{b}(x)^{2}\leq 2(a)^{2}+2(\bar{\Gamma}-1)^{2}(b)^{2}$, where
@@
[…]
@@
by the triangle inequality and using $u_{0,0}(x),u_{0,1}(x),u_{1,0}(x),u_{1,1}(x)$ are non-negative and sum to one. Substituting back into the bound on $R(\hat{d})$ delivers
[…]

### v6

@@ 5.1 Bounding Individual Default Risk
We construct robust credit risk scores using our estimator from Section 3, exploring how varying confounding assumptions affect individual default risk predictions. We estimate the upper bound on the conditional default probability with observed outcome bounds, $\overline{\mu}(\cdot)$. We split applications into two folds: the first estimates $\pi_{1}(\cdot)$, $\mu_{1}(\cdot)$ using random forests; the second regresses estimated pseudo-outcomes on application-level covariates $X_{i}$ using cross-validated Lasso regression. We set $\underline{\Gamma}=1$ and vary $\overline{\Gamma}\in\{1,2,3\}$.
We compare our robust estimator (with $\underline{\Gamma}=1,\overline{\Gamma}=2$) against a benchmark score trained only on funded applications. Figure 1’s left panel shows the joint distribution of their predictions. For each benchmark decile, the heatmap displays the percentage of applications falling into each decile of our robust predictions. Applications exhibit substantial reranking: among those in the benchmark’s 5th decile, 18.7% shift to extreme deciles (1-3 or 8-10) under the robust score. These differences meaningfully affect credit access. For instance, if only the least risky third of applications receive funding, 10.1% would have their decision reversed when comparing benchmark versus robust scores.
Figure 1’s right panel examines how estimated coefficients vary with $\overline{\Gamma}$ (variable definitions in Online Appendix Table A3). Confounding assumptions affect different covariates heterogeneously. Some coefficients—total net income and credit bureau score—remain zero across all $\overline{\Gamma}$ values. Others, like the number of recent credit card applications, receive zero weight in the benchmark but non-zero weight for all $\overline{\Gamma}$¿1. Still others, such as occupation type and maximum recent delinquency, exhibit large magnitude changes as $\overline{\Gamma}$ varies, highlighting how confounding assumptions shape which applicant characteristics the algorithm prioritizes.
5.2 Bounding the Predictive Performance of a Credit Risk Score
We now evaluate an existing credit risk score’s performance. We split applications into training and evaluation data. On the training data, we construct a benchmark score predicting 5-month default risk among funded applications using random forests. On the evaluation data, we analyze mean square error and ROC curves, following recent mortgage lending studies (BlattnerNelson(21); FusterEtAl(22)predictably). Using observed outcome bounds with $\underline{\Gamma}=1$, we assess how the benchmark’s estimated performance varies as we increase $\overline{\Gamma}$, allowing unfunded applicants to have progressively higher default rates than funded applicants with similar characteristics.
Figure 2(a) shows how estimated MSE bounds vary with $\overline{\Gamma}$. When $\overline{\Gamma}=1$ (the unconfounded case), both bounds equal the benchmark’s MSE on funded applications (dashed red line; $0.143$). As $\overline{\Gamma}$ increases, allowing unfunded applicants higher default rates, bounds widen but remain informative. At $\overline{\Gamma}=2$ — permitting unfunded applicants to default at twice the rate of funded applicants with similar characteristics — the estimated bounds are $[0.138,0.182]$. Figure 2(b) examines how estimated ROC curves vary with $\overline{\Gamma}$. We summarize ROC curves using Area Under the Curve (AUC), computed via the trapezoidal rule. Under unconfoundedness ($\overline{\Gamma}=1$), the benchmark’s AUC is $0.637$. At $\overline{\Gamma}=1.5$, AUC bounds are $[0.575,0.705]$. As with MSE, bounds widen with $\overline{\Gamma}$ but remain informative at empirically plausible confounding levels.
We finally investigate the benchmark credit risk score’s predictive disparities across income groups. We define $G_{i}\in\{0,1\}$ to be whether an application is below or above the median personal income of all submitted applications. We examine how the benchmark credit risk score’s mean square error and ROC curve vary across these income groups, thereby studying whether it is a “noisier” signal of 5-month default risk for lower income applications, as in BlattnerNelson(21); FusterEtAl(22)predictably.
Figure 3(a) compares the mean square error of the benchmark credit risk across income groups. Interestingly, for $\overline{\Gamma}=1$, the mean square error of the benchmark credit risk score is significantly larger on applications below the median income (0.169) than those above it (0.136). This difference persists as we vary assumptions on unobserved confounding. The upper bound on the mean square error for applications above the median income is only larger than the lower bound on the mean square error for applications below the median income for values $\overline{\Gamma}\geq 1.69$. Ruling out this mean square error disparity across income groups would require that unfunded applications are at least 1.69 times as likely to default within 5 months as funded applications conditional on covariates. [^8] Rather than examining disparity bounds across a range of $\overline{\Gamma}$ values, researchers could compute the “identification breakdown point” $\overline{\Gamma}^{*}$: the minimal value of $\overline{\Gamma}$ (holding $\underline{\Gamma}=1$) such that $0\in H(\text{disp}(s;\beta))$. A natural estimator is $\widehat{\overline{\Gamma}}^{*}$, the minimal $\overline{\Gamma}$ such that zero is contained in the estimated $(1-\alpha)$ confidence set $\widehat{H}(\text{disp}(s;\beta))$. Since our confidence sets have valid coverage by the results in Section 4.1 and Appendix C.1, $P(\widehat{\overline{\Gamma}}^{*}\geq\overline{\Gamma}^{*})\geq 1-\alpha$. Figure 3(b) also compares the ROC curves of the benchmark credit risk score across income groups. Intriguingly, there exists effectively no disparity across income groups based on the ROC curves. When $\overline{\Gamma}=1$, the implied AUC for applications above the median income is $0.603$, whereas the implied AUC is $0.616$ for applications below the median income. Any disparity in terms of ROC curves is already ruled out at $\overline{\Gamma}=1.25$ as the bounds on the AUC for applications above the median income are $[0.567,0.653]$ and $[0.581,0.654]$ for applications below the median income.
6 Conclusion
We developed a framework for robust design and evaluation of predictive algorithms under unobserved confounding. Our approach nests popular empirical strategies — observed outcome bounds, proxy outcomes, and instrumental variables — by bounding outcome differences between selected and unselected units (Assumption 2.1). We provide estimators that work well across bounding strategies and performance measures: conditional likelihoods for algorithm design, overall metrics like MSE and calibration, and class-specific measures like TPR and FPR. This breadth makes our framework valuable across domains where predictive algorithms face selective labels—pretrial release, consumer lending, medical diagnosis, hiring, and child welfare.
Several open questions present promising research directions. First, achieving full double robustness for positive class performance bounds remains open, though our technical results suggest possible paths forward. Second, uniform inference for breakdown frontiers would let researchers formally test how strong confounding must be to overturn conclusions (MastenPoirier(20)). Third, we evaluate fixed algorithms; extending our results to settings where the same data is used to train and evaluate an algorithm is an important direction. Finally, extending to multi-valued or continuous outcomes would broaden applicability, though defining class-specific performance in these settings requires care.
As algorithms increasingly drive high-stakes decisions in credit, criminal justice, and healthcare, robust evaluation under realistic confounding becomes essential. Our framework and these open questions offer pathways to strengthen algorithmic accountability where the stakes are highest.
Main Figures
@@ 6 Conclusion
Robust Design and Evaluation of Predictive Algorithms
under Unobserved Confounding
Online Appendix
Ashesh Rambachan     Amanda Coston     Edward H. Kennedy
This online appendix contains additional results for “Robust Design and Evaluation of Predictive Algorithms under Unobserved Confounding” by Ashesh Rambachan, Amanda Coston, and Edward Kennedy. Section A contains an oracle inequality for the integrated mean square convergence rate of the pseudo-outcome regression procedures. Section B contains the proofs of all theoretical results presented in the main text. Section C contains additional theoretical results discussed in the main text. Section D extends our estimation results to alternative bounding functions, such as proxy outcome bounds and instrumental variable bounds. Section E discusses connections between existing sensitivity analysis frameworks in causal inference and observed outcome bounds in our framework. Section F contains Monte Carlo simulations, and Section G contains additional empirical results.
Appendix A Oracle Inequality for Pseudo-Outcome Regression
We provide an oracle inequality on the $L_{2}(\mathbb{P})$-error of nonparametric regression with estimated pseudo-outcomes. This generalizes the pointwise analysis in Kennedy(22)-towardsoptimal. In Section 3 of the main text, we apply this oracle inequality to analyze our estimators for the conditional probability bounds.
We first state a $L_{2}(\mathbb{P})$-stability condition on the second-stage nonparametric regression estimator. We then show that the $L_{2}(\mathbb{P})$-stability condition is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest model and several others.
Assumption A.1.
$\frac{\int\Big[\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})\mid X_{i}=x\}-\widehat{\mathbb{E}}_{n}\{{f}(V_{i})\mid X_{i}=x\}-\widehat{\mathbb{E}}_{n}\{\widehat{b}(X_{i})\mid X_{i}=x\}\Big]^{2}\ d\mathbb{P}(x)}{\mathbb{E}\left(\int\Big[\widehat{\mathbb{E}}_{n}\{f(V_{i})\mid X_{i}=x\}-\mathbb{E}\{f(V_{i})\mid X_{i}=x\}\Big]^{2}d\mathbb{P}(x)\right)}\ \xrightarrow{p}0$ (8)
whenever $d(\widehat{f},f)\xrightarrow{p}0$.
Proposition A.1.
$d(\widehat{f},f)=\|\widehat{f}-f\|_{w^{2}}\equiv\sum_{i=1}^{n}\left\{\frac{\|w_{i}(\cdot;X^{n})\|^{2}}{\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}}\right\}\int\left\{\widehat{f}(v)-f(v)\right\}^{2}\ d\mathbb{P}(v\mid X_{i}),$
Proof.
and this term has mean zero since
by definition of $\widehat{b}$ and iterated expectation. Therefore,
$\leq\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\int\Big\{\widehat{f}(v)-f(v)\Big\}^{2}\ d\mathbb{P}(v\mid X_{i})$
$=\|\widehat{f}-f\|_{w^{2}}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}$
Further note that $R_{n}^{2}$ equals
$\mathbb{E}$ $[\|\widetilde{m}-m\|]^{2}=\mathbb{E}\left(\int\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{f(V_{i})-m(X_{i})\Big\}+\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right]^{2}d\mathbb{P}(x)\right)$
$=\mathbb{E}\left(\int\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{f(V_{i})-m(X_{i})\Big\}\right]^{2}d\mathbb{P}(x)\right)+\mathbb{E}\left[\int\left\{\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right\}^{2}d\mathbb{P}(x)\right]$
$=\mathbb{E}\left\{\int\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ \sigma(X_{i})^{2}\ d\mathbb{P}(x)\right\}+\mathbb{E}\left[\int\left\{\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right\}^{2}d\mathbb{P}(x)\right]$
@@ Appendix A Oracle Inequality for Pseudo-Outcome Regression
Lemma A.1.
Under the same setup from Assumptions A.1, define (i) $m(x)=\mathbb{E}[f(V_{i})\mid X_{i}=x]$ the conditional expectation of $f(V_{i})$ given $X_{i}$; (ii) $\hat{m}(x):=\hat{\mathbb{E}}_{n}[\hat{f}(V_{i})\mid X_{i}=x]$ the regression of $\hat{f}(V_{i})$ on $X_{i}$ in the test samples; (iii) $\tilde{m}(x):=\hat{\mathbb{E}}_{n}[f(V_{i})\mid X_{i}=x]$ the oracle regression of $f(V_{i})$ on $X_{i}$ in the test samples. Furthermore, let $\tilde{b}(x):=\hat{\mathbb{E}}_{n}[b(V_{i})\mid X_{i}=x]$ be the $\hat{\mathbb{E}}_{n}$-smoothed bias and $R_{n}^{2}=E[\|\widetilde{m}-m\|]^{2}$ be the oracle $L_{2}$-error. If
  1. i.
  2. ii.
    $d(\hat{f},f)\xrightarrow{p}0$,
then
$\|\hat{m}-\tilde{m}\|=\|\tilde{b}(\cdot)\|+o_{\mathbb{P}}(R_{n}).$
$\frac{\|\widehat{m}-\widetilde{m}\|}{\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}}\ \xrightarrow{p}\ 0$
and
$\|\widehat{m}-m\|=\|\tilde{m}-m\|+o_{\mathbb{P}}(R_{n}).$
Proof.
Note
This generalizes Proposition 1 of Kennedy(22)-towardsoptimal, which shows that a pointwise stability condition and consistency of $\hat{f}$ implies an oracle inequality on the pointwise convergence of a feasible pseudo-outcome regression. In Section 3 of the main text, we apply Lemma A.1 to analyze the convergence of our proposed estimators for the conditional probability bounds.
Appendix B Proofs of Results in the Main Text
B.1 Proof of Lemma 2.1
@@ Appendix B Proofs of Results in the Main Text
[…]
Towards this, we apply a change-of-variables. Let $U(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ be defined as $U(x)=\frac{\delta(x)-\underline{\delta}(x)}{\overline{\delta}(x)-\underline{\delta}(x)}$. For any $\delta(\cdot)\in\Delta$, there exists $U(\cdot)\in[0,1]$ such that $\mbox{perf}_{+}(s;\beta,\delta)=\mbox{perf}_{+}(s;\beta,U)$, where
@@
[…]
@@
[…]
B.2 Proof of Lemma 2.2
Define the bounds
@@
This is minimized over $\mu^{*}(x)\in\mathcal{H}(\mu^{*}(x))$ at $\mu^{*}(x)=\underline{\mu}^{*}(x)$. The result for the lower bound immediately follows. The upper bound follows analogously. $\Box$
B.3 Proof of Proposition 3.1
We prove the result for our estimator of the upper bound, and the same argument applies for our estimator of the lower bound. We define $\widehat{\pi}_{0}(x)=1-\widehat{\pi}_{1}(x)$ throughout. Let $\mathcal{O}_{1}$ denote the observations in the first fold and $\mathcal{O}_{2}$ denote the observations in the second fold. We first observe that
@@
Under Assumption A.1, $\|\widehat{\overline{\mu}}(\cdot)-\widehat{\overline{\mu}}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma A.1. Furthermore, $\hat{b}(x)^{2}\leq 2(a)^{2}+2(\bar{\Gamma}-1)^{2}(b)^{2}$, where
@@
$R(\hat{d})=\int_{x\in\mathcal{X}}\left(c(x)-\tilde{\mu}^{*}(x)\right)\left(d^{*}(x)-\hat{d}(x)\right)dP(x)\leq\int_{x\in\mathcal{X}}\left|\tilde{\mu}^{*}(x)-c(x)\right|1\{d^{*}(x)\neq\hat{d}(x)\}dP(x).$
At any fixed $X_{i}=x$, $\hat{d}(X_{i})\neq d^{*}(X_{i})$ implies that $|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|\geq|\tilde{\mu}^{*}(x)-c(x)|$. Combining this with the previous display implies $R(\hat{d})\leq\int_{x\in\mathcal{X}}|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|dP(x).$ Substituting in the definition of $\tilde{\mu}^{*}(x)$ and $\widehat{\tilde{\mu}}(x)$, we have
@@
by the triangle inequality and using $u_{0,0}(x),u_{0,1}(x),u_{1,0}(x),u_{1,1}(x)$ are non-negative and sum to one. Substituting back into the bound on R(d^)R($\hat${d}) delivers
$R(\hat{d})\leq\int_{x\in\mathcal{X}}|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)|dP(x)+\int_{x\in\mathcal{X}}|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)|dP(x)=\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{1}+\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{1}.$